OpenStudy (anonymous):

sole algebraically. (x-1)^2=16

OpenStudy (saifoo.khan):

take squareroot on both sides.

OpenStudy (anonymous):

x-1=4?

OpenStudy (saifoo.khan):

\[x -1 = \sqrt{16}\] \[x-1 = 4\] \[x = 4+1\] \[x = 5\]

jhonyy9 (jhonyy9):

- so and -3 is right too

jhonyy9 (jhonyy9):

because -3-1=-4 squared =16